### v1

# Gaussian Scenes: Pose-Free Sparse-View Scene Reconstruction using Depth-Enhanced Diffusion Priors

Abstract
In this work, we introduce a generative approach for pose-free reconstruction of $360^{\circ}$ scenes from a limited number of uncalibrated 2D images. Pose-free scene reconstruction from incomplete, unposed observations is usually regularized with depth estimation or 3D foundational priors. While recent advances have enabled sparse-view reconstruction of unbounded scenes with known camera poses using diffusion priors, these methods rely on explicit camera embeddings for extrapolating unobserved regions. This reliance limits their application in pose-free settings, where view-specific data is only implicitly available. To address this, we propose an instruction-following RGBD diffusion model designed to inpaint missing details and remove artifacts in novel view renders and depth maps of a 3D scene. We also propose a novel confidence measure for Gaussian representations to allow for better detection of these artifacts. By progressively integrating these novel views in a Gaussian-SLAM-inspired process, we achieve a multi-view-consistent Gaussian representation. Evaluations on the MipNeRF360 dataset demonstrate that our method surpasses existing pose-free techniques and performs competitively with state-of-the-art posed reconstruction methods in complex $360^{\circ}$ scenes.
1 Introduction
Obtaining high-quality 3D reconstructions or novel views from a sparse set of images has been a long-standing goal in computer vision. Recent methods for sparse-view reconstruction often employ generative, geometric, or semantic priors to stabilize the optimization of NeRFs (Mildenhall et al., 2020) or Gaussian splats (Kerbl et al., 2023) in highly under-constrained scenarios. However, they typically assume access to accurate intrinsic and extrinsic parameters, often derived from dense observations. This reliance on ground-truth poses is a restrictive assumption, making these methods impractical for real-world applications. Moreover, pose estimation from sparse views is error-prone; both traditional Structure from Motion approaches and recent 3D foundational models (Wang et al., 2024; Leroy et al., 2024) struggle with insufficient matching features between image pairs. In response, recent pose-free approaches using 3D Gaussian splatting (3DGS) integrate monocular depth estimation (Ranftl et al., 2020), 2D semantic segmentation (Kirillov et al., 2023), or 3D foundational priors (Wang et al., 2024), optimizing 3D Gaussians and camera poses together during training. However, these methods are typically designed for scenes with high view overlap, and they often fail to reconstruct complex, large-scale $360^{\circ}$ scenes with sparse coverage. Additionally, despite extensive regularization to prevent overfitting, the limited observations impede coherent synthesis of unobserved regions. This challenge underlines the need for additional generative regularization to enable accurate extrapolation and complete scene reconstruction.
Despite several previous attempts (Deng et al., 2022; Roessle et al., 2022; Li et al., 2024; Xiong et al., 2023; Zhu et al., 2023; Chung et al., 2023; Wang et al., 2023a; Jain et al., 2021; Niemeyer et al., 2022; Wynn and Turmukhambetov, 2023) in the field of sparse novel view synthesis, only a few - CF-3DGS (Fu et al., 2023), COGS Jiang et al. (2024), InstantSplat (Fan et al., 2024) have been successful in dealing with the problem without precomputed camera parameters. These methods jointly optimize Gaussians and camera parameters during training and use different 2D priors for regularization. Although they find success in artifact and blur reduction, they are less effective in $360^{\circ}$ scene reconstruction due to a lack of generative priors.
Reconstructing full $360^{\circ}$ scenes from limited views requires robust priors, such as those from powerful 2D image or video diffusion models (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022) that encode common 3D structures. Recent methods (Liu et al., 2023; Sargent et al., 2024; Wu et al., 2024; Gao* et al., 2024; Blattmann et al., 2023) incorporate view and context conditioning into these models, fine-tuning them on extensive real-world and synthetic multi-view datasets to support radiance field optimization. This enables realistic extrapolation into unobserved areas of complex $360^{\circ}$ scenes. However, these models depend on accurate pose information, making them unsuitable for scenarios where pose estimates deviate significantly from ground truth, leading to misaligned view generation. To our knowledge, only Gaussian Object (Yang et al., 2024), iFusion (Wu et al., 2023), and UpFusion (Nagoor Kani et al., 2024) provide pose-free generative solutions, but they are designed specifically for 3D object reconstruction, not large-scale scenes.
In the absence of camera conditioning, we augment Stable Diffusion Rombach et al. (2022) model with additional channels for context, 3DGS render, and depth maps with Gaussian artifacts, and a pixel-aligned confidence map capturing missing regions and reconstruction artifacts in the RGBD image. During inference, the model predicts a clean, inpainted version of the conditioning artifact image and an aligned depth map. This formulation prevents a pose-conditioning requirement common with most multiview diffusion models while also using limited multiview data for finetuning. This also alleviates dependency on large-scale 3D datasets which are usually synthetic and low quality compared to real-world scenes.
We present GScenes, an efficient method that uses 3D foundational (Leroy et al., 2024) and RGBD diffusion priors for pose-free sparse-view reconstruction of complex $360^{\circ}$ scenes. We first estimate a point cloud and approximate scene intrinsics and extrinsics using DUSt3R and then jointly optimize both Gaussians and camera parameters with just a few iterations of 3DGS. We then generate novel views by sampling from a B-spline trajectory fitted to the training views. Resulting renders and depth maps contain missing details and Gaussian artifacts that are refined by our diffusion prior and then used to further optimize the 3D Gaussians and camera poses. Despite using weaker generative priors compared to state-of-the-art pose-dependent methods, we show competitive qualitative and quantitative performance with ReconFusion (Wu et al., 2024) and CAT3D (Gao* et al., 2024) and comprehensively outperform all other pose-dependent and pose-free techniques on reconstruction quality. Our method leverages stronger priors than simple regularizers while not relying on million-scale multi-view data or huge compute resources to train a 3D-aware diffusion model.
In summary, our contributions are as follows.
  • We propose a confidence measure to effectively detect empty regions, floaters, and artifacts in novel view renders and depth maps from 3DGS.
  • We present an instruction-following RGBD diffusion model capable of synthesizing plausible novel views in a 3DGS scene initialized from sparse unposed images.
  • We show that integrating our diffusion priors with a strong geometry prior in DUSt3r facilitates fast reconstruction of a 3D scene previously only possible with pose-dependent finetuned video diffusion priors.
  • We show that GScenes outperforms recent works based on regularization and generative priors in the reconstruction of large 3D scenes.
2 Related Work
2.1 Sparse-View Radiance Fields
After NeRF Mildenhall et al. (2020) and 3D Gaussian Splatting Kerbl et al. (2023) advanced inverse rendering of radiance fields, numerous methods emerged to reduce dense scene capture needs to sparse views, categorized into regularization techniques and generalizable reconstruction priors.
Regularization Techniques
Fitting a 3D representation from sparse observations is an ill-posed problem prone to local minima, often manifesting as ’floaters’ in novel view renderings of radiance fields. Regularization techniques are commonly used to address limited data. Existing methods leverage depth from Structure-from-Motion (SfM)Deng et al. (2022); Roessle et al. (2022), monocular estimationLi et al. (2024); Xiong et al. (2023); Zhu et al. (2023); Chung et al. (2023), or RGB-D sensors Wang et al. (2023a). DietNeRF Jain et al. (2021) uses a semantic consistency loss based on CLIP Radford et al. (2021) features, while FreeNeRF Yang et al. (2023) regularizes the frequency range of NeRF inputs. Approaching generative priors, RegNeRF Niemeyer et al. (2022) and DiffusioNeRF Wynn and Turmukhambetov (2023) maximize the likelihoods of rendered patches using normalizing flows or diffusion models, respectively.
Generalizable Reconstruction
When only a few or a single view is available, regularization techniques are often insufficient to resolve reconstruction ambiguities. To address this, recent research focuses on training priors for novel view synthesis across multiple scenes. pixelNeRF Yu et al. (2021) uses pixel-aligned CNN features as conditioning for a shared NeRF MLP, while other approaches Trevithick and Yang (2021); Chen et al. (2021); Henzler et al. (2021); Lin et al. (2023b) condition NeRF on 2D or fused 3D features. Further priors have been learned on triplanes Irshad et al. (2023), voxel grids Guo et al. (2022), and neural points Wewer et al. (2023). Leveraging 3D Gaussian Splatting Kerbl et al. (2023), methods like pixelSplat Charatan et al. (2024) and MVSplat Chen et al. (2024) achieve state-of-the-art performance in stereo view interpolation and real-time rendering. However, these regression-based models often produce blurry outputs under high uncertainty. In contrast, generative methods like GeNVS Chan et al. (2023) and latentSplat Wewer et al. (2024) aim to sample from multi-modal distributions, offering better handling of ambiguous novel views.
Generative Priors
For ambiguous novel views, predicting expectations over all reconstructions may be unreliable. Consequently, regression approaches fall short, whereas generative methods attempt to sample from a multi-modal distribution.
Diffusion Models
In recent years, diffusion models Ho et al. (2020); Dhariwal and Nichol (2021) emerged as the state-of-the-art for image synthesis. They are characterized by a pre-defined forward noising process that gradually destroys data by adding random (typically Gaussian) noise. The objective is to learn a reverse denoising process with a neural network that, after training, can sample from the data distribution given pure noise. Important improvements include refined sampling procedures Song et al. (2021); Karras et al. (2022) and the more efficient application in a spatially compressed latent space compared to the high-resolution pixel space Rombach et al. (2022). Their stable optimization, in contrast to GANs, enabled today’s text-to-image generators Nichol et al. (2022); Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022) trained on billions of images Schuhmann et al. (2022). In recent years, diffusion models Ho et al. (2020); Dhariwal and Nichol (2021) have become state-of-the-art for image synthesis. These models use a forward noising process to progressively add random (typically Gaussian) noise, and train a neural network to reverse this process, enabling sampling from the data distribution given pure noise. Key improvements include refined sampling techniques Song et al. (2021); Karras et al. (2022) and more efficient application in compressed latent spaces rather than high-resolution pixel spaces Rombach et al. (2022). Their stable optimization, unlike GANs, has enabled the success of text-to-image generators Nichol et al. (2022); Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022) trained on billions of images Schuhmann et al. (2022).
2D Diffusion for 3D While diffusion models have been applied directly on 3D representations like triplanes Shue et al. (2023); Chen et al. (2023a), voxel grids Müller et al. (2023), or (neural) point clouds Zhou et al. (2021); Melas-Kyriazi et al. (2023); Schröppel et al. (2024), 3D data is scarce. Given the success of large-scale diffusion models for image synthesis, there is a great research interest in leveraging them as priors for 3D reconstruction and generation. DreamFusion Poole et al. (2023) and follow-ups Wang et al. (2023b); Lin et al. (2023a); Chen et al. (2023b); Deng et al. (2023); Tang et al. (2024) employ score distillation sampling (SDS) to iteratively maximize the likelihood of radiance field renderings under a conditional 2D diffusion prior. For sparse-view reconstruction, existing approaches incorporate view-conditioning via epipolar feature transform Zhou and Tulsiani (2023), cross-attention to encoded relative poses Liu et al. (2023); Sargent et al. (2024), or pixelNeRF Yu et al. (2021) feature renderings Wu et al. (2024). However, this fine-tuning is expensive and requires large-scale multi-view data, which we circumvent with GScenes.
Figure 1: Overview of GScenes. We render 3D Gaussians fitted to our sparse set of $M$ views from a novel viewpoint. The resulting render and depth map have missing regions and Gaussian artifacts, which are rectified by an RGBD image-to-image diffusion model. This then acts as pseudo ground truth to spawn and update 3D Gaussians and satisfy the new view constraints. This process is repeated for several novel views spanning the $360^{\circ}$ scene until the representation becomes multi-view consistent.
3 Method
This section begins with an overview of our pipeline in Sec 3.1. In Sec 3.3, we present a confidence measure for 3DGS rasterization to capture artifacts and missing regions in rendered images and depth maps. We detail the different components of our generative prior in Sec 3.2 and the subsequent 3DGS optimization in Sec 3.4. We end the section with a final pose alignment step for a given set of test views in Sec 3.5.
3.1 Algorithm Overview
Problem Setup
Given a set of $M$ images =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{M}\}$ of an underlying 3D scene with unknown intrinsics and extrinsics, our goal is to reconstruct the 3D scene, estimate the camera poses of a monocular camera at the $M$ training views, and synthesize novel views at evaluation time given by $N$ unseen test images $\{I_{M+1},I_{M+2},\dots,I_{M+N}\}$.
Algorithm 1 GScenes Algorithm
1:Sparse input image set =$\mathcal{I}=\{I_{1},I_{2},\dots I_{M}\}$
2:Set of 3D Gaussians $\mathcal{G}$ and camera poses $\pi=\{\pi_{1},\pi_{2},\dots\pi_{M}\}$
$\hat{\mathcal{I}}\leftarrow\mathcal{I}$
4:$\mathcal{G}\leftarrow$ Optimize 3DGS with DUSt3r point cloud for $1k$ iterations
5:for $N$ iterations do
6:     $\pi\leftarrow$ Sample novel camera pose.
7:     $\mathbf{I},\mathbf{D}\leftarrow R_{\pi}(\mathcal{G})$ Render from camera $\pi$
8:     $\mathbf{I}\leftarrow$ Refine$(\mathbf{I})$
9:     $\hat{\mathbf{I}}\leftarrow\ \hat{\mathbf{I}}\cup\{\mathbf{I}\}$
10:     $\mathcal{G}\leftarrow$ Optimize 3DGS for $k$ iterations
11:end for
An overview of our method is given in Fig 1 and Alg 1. We initialize GScenes with an incomplete dense Gaussian point cloud reconstruction from sparse input images using InstantSplat Fan et al. (2024). We use this incomplete Gaussian scene representation as an implicit geometric prior and sample novel views along a smooth B-spline trajectory fitted to training views. We use our customized RGBD diffusion prior to generate plausible novel views. In addition to CLIP features of source images for context, we devise a novel 3DGS confidence measure to effectively guide our diffusion model towards empty regions and potential artifacts in a novel view render. We then run $10k$ iterations of 3DGS optimization, sampling a novel view before each densification step (Fu et al., 2023) to obtain our final scene representation.
Gaussian Point Cloud Initialization
The DUSt3r pipeline gives us a pixel-aligned dense-stereo point cloud $\mathbf{P}\in\mathbb{R}^{S\times 3}$, camera intrinsics $\{\mathbf{K_{i}}\in\mathbb{R}^{3\times 3}\}_{i=1}^{M}$ and extrinsics $\{\mathbf{E_{i}}=[\mathbf{R_{i}|T_{i}}]\}_{i=1}^{M}$ for our $M$ input images. Nevertheless, both $\mathbf{P}$ and the estimated poses demonstrate sub-optimal alignment compared to those generated by COLMAP (Schönberger and Frahm, 2016) from the dense observation dataset. Consequently, similar to the approach used in InstantSplat, we start by initializing 3D Gaussians at each location within the globally aligned point cloud $\mathbf{P}$. Subsequently, we jointly optimize both the Gaussian attributes and camera parameters over the course of $1k$ iterations without incorporating any form of Adaptive Density Control.
3.2 RGBD Diffusion Priors for NVS
Despite the initial geometric prior, many areas of the scene remain unobserved due to the sparsity of the input set, leading to areas of no Gaussians (0-Gaussians) in the 3D representation $\mathcal{G}$ obtained from InstantSplat and zero opacity regions, floaters, and artifacts in novel views. Regularization techniques or 3D foundational priors cannot help with inferring missing details, prompting us to develop an instruction-following RGBD diffusion model comprising a variational autonecoder with encoder $\mathcal{E}$, decoder $\mathcal{D}$ and UNet $\bm{\epsilon_{\theta}}$ operating in the latent space of the VAE. $\bm{\epsilon_{\theta}}$ learns to predict noise in a noisy latent $z_{t}$ conditioned on a RGBD image $\hat{I}$, corresponding confidence map $\mathcal{C}$, clip conditioning $c_{clip}$, and a text instruction $y$. $\hat{I}$ and $\mathcal{C}$ are obtained by rendering $\mathcal{G}$ from novel views.
We initialize $\bm{\epsilon_{\theta}}$ with the pretrained weights of the Stable-Diffusion-2 checkpoint(Rombach et al., 2022) to leverage its image generation capabilities for RGBD image-to-image synthesis. We introduce additional channels in the first convolutional layer of $\bm{\epsilon_{\theta}}$ to support image conditioning with $\hat{I}$. Specifically, we concatenate $z_{t}$, $\mathcal{E}(\hat{I})$ and $\mathcal{\hat{C}}$, where $\mathcal{\hat{C}}$ is $\mathcal{C}$ downsampled to the resolution of the latent space. Weights of all newly added channels are initialized to zero. However, conditioning on $\hat{I}$ and $\mathcal{C}$ alone does not ensure novel view synthesis coherent with the observed inputs. In the absence of explicit 3D conditioning, we augment our generative model with scene context using CLIP features. We extract CLIP features $c_{clip}\in\mathcal{R}^{M\times d}$ of the $M$ training images to modulate intermediate feature layers in $\bm{\epsilon_{\theta}}$ However, instead of the usual cross-attention mechanism, we resort to a more parameter-efficient approach by using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018). These layers work by generating scaling and shifting parameters for the UNet feature maps, allowing $c_{clip}$ to modulate intermediate activations in the UNet. We first pass $c_{clip}$ through trainable self-attention modules to capture shared semantic information across views and create a unified, context-aware representation.
$c_{attn}=\text{SelfAttention}(c_{clip})$ (1)
$c_{attn}\in\mathcal{R}^{d}$ is next passed through layer-specific fully connected layers to predict scale and shift parameters $\gamma,\beta$ as:
$\gamma^{(l)},\beta^{(l)}=FC^{(l)}(c_{attn})$ (2)
Let $\mathbf{F}^{(l)}\in\mathcal{R}^{k\times h\times w}$ represent a feature map in $\bm{\epsilon_{\theta}}$ at a layer $l$, where $k$ is the number of channels, $h$ and $w$ are the spatial dimensions. The FiLM modulation is applied element-wise as:
$\mathbf{F}_{mod}^{(l)}=\text{FiLM}(\mathbf{F}^{(l)}|\gamma^{(l)},\beta^{(l)})=\gamma^{(l)}\cdot\mathbf{F}^{(l)}+\beta^{(l)}$ (3)
This conditioning enables the dynamic adaptation of the UNet’s feature maps facilitated by scale and shift parameters conditioned on the CLIP embedding, thereby directing the UNet’s output to align with the semantic features captured by CLIP. We implement FiLM modulation with CLIP embeddings exclusively within the down and mid blocks of the UNet, as this approach allows for the effective integration of conditional information without perturbing the generation process in the up(sample) blocks.
3.3 Pixel-Aligned Confidence Map
To effectively guide our diffusion model towards recognizing and repairing artifacts in novel views, we design a novel pixel-aligned confidence measure based on the cumulative transmittance along a viewing direction at a pixel $i$ and the number of overlapping Gaussians at that pixel. Formally, we define confidence at a pixel as:
$\mathcal{C}_{i}=-\log(T_{i}+\epsilon)\times n_{contrib}$ (4)
where $T_{i}=\prod_{i}(1-\alpha_{i})$ represents the fraction of light transmitted through pixel $i$ without interaction and $n_{contrib}$ is the number of contributing 3D Gaussians. $\epsilon>0$ is a small constant to prevent taking the logarithm of 0. Intuitively, low $T_{i}$ means the viewing ray interacts significantly with 3D Gaussians, implying higher confidence in the pixel’s color. The negative logarithm provides a more perceptually uniform confidence scale. $n_{contrib}$ further increases (or decreases) the reliability of the rendered color or depth due to consensus among multiple (few) Gaussian primitives.
To the best of our knowledge, only 3DGS-Enhancer (Liu et al., 2024) proposes a confidence measure based on alpha-compositing of Gaussian scales, assuming that well-reconstructed areas are represented by Gaussians of small volumes, whereas large scales usually point towards artifacts. However, as can be seen in Fig 2, this assumption does not hold true in the case of regions with monotonous textures, which do not necessarily need multiple Gaussians of small scales to represent fine details. One can clearly see that our proposed measure effectively captures both empty regions as well as reconstruction artifacts, avoiding false positives unlike  Liu et al. (2024).
Render Conf. Measure  (Liu et al., 2024) Conf. Measure (Ours)
DTU scan38 (6) Refer to caption Refer to caption Refer to caption
DTU scan25 (6) Refer to caption Refer to caption Refer to caption
Figure 2: Confidence Measure comparison with 3DGS-Enhancer (Liu et al., 2024). Our confidence map accurately identifies artifacts and 00-Gaussian regions in the image (darker pixels) while  (Liu et al., 2024) incorrectly attributes high confidence to regions with overlap of small-scale Gaussians.
3.3.1 Synthetic RGBD Dataset Creation
Figure 3: Training our RGBD diffusion model. Pairs of RGBD clean images and images with artifacts are obtained from 3DGS fitted to sparse and dense observations, respectively, across $165$ scenes. CLIP features provide scene context, and a confidence map additionally detects empty regions and artifacts in the artifact image. These are combined with one of 51 synonymous prompts generated by GPT-4 Achiam et al. (2023) from a base instruction. The Stable Diffusion UNet is then fine-tuned with a dataset of $28,161$ samples.
For training the additional weights in $\bm{\epsilon_{\theta}}$ for RGBD image-to-image diffusion, we rely on a set $\mathcal{X}=\{(I^{i},\hat{I}^{i},\mathcal{C}^{i},y^{i})_{i=1}^{N}\}$ of data quartets, each containing a clean RGBD image $I^{i}$, an RGBD image with artifacts $\hat{I}^{i}$ and the corresponding confidence map $\mathcal{C}^{i}$, and an edit instruction for fine-tuning $y^{i}$, to “teach” the diffusion model how to detect Gaussian artifacts guided by the confidence map and generate a clean version of the conditioning image. For this, we build an artifact simulation engine comprising a high-quality 3DGS model fitted to dense views, a low-quality 3DGS model fitted to few views, and camera interpolation and perturbation modules to use supervision of the high-quality model at viewpoints beyond ground truth camera poses. The fine-tuning setup is illustrated in Fig. 3. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. To have diversity in edit instructions, we start with a base human-intuitive instruction - “Inpaint low-confidence regions in the image and depth map with coherent details and remove all floaters and Gaussian artifacts.” and ask GPT-4 to generate 50 synonymous instructions. During training, each image triplet is randomly combined with one of these 51 instructions. Additionally, for each $M$, we save the CLIP features $c_{clip}$ of the $M$ sparse views for conditioning $\bm{\epsilon_{\theta}}$
3.3.2 Depth-Augmented Autoencoder Finetuning
For encoding and decoding RGBD images, we customize a Variational AutoEncoder by introducing additional channels in the first and last convolutional layers of the Stable Diffusion VAE. A similar approach was followed by Stan et al. (2023), where their KL-autencoder was finetuned with triplets containing RGB images, depth maps, and captions to train the weights in the new channels. However, the depth maps used for fine-tuning this VAE were estimated using MiDaS(Ranftl et al., 2020), which are usually blurry monocular depth estimates. As such, reconstructing RGBD images using this VAE produces depth maps with extreme blur - not ideal for a scene reconstruction problem. Hence, we further finetune this VAE with our synthetic dataset, which contains depth maps rendered by the differentiable 3DGS rasterizer, giving accurate pixel depth with high-frequency details. Specifically, we use the following objective:
$\mathcal{L}_{\text{autoencoder}}=$ $\underset{\mathcal{E},D}{\text{min}}\,\underset{D_{\psi}}{\text{max}}\big{(}\mathcal{L}_{\text{rec}}(x,D(\mathcal{E}(x)))$ (5)
$-\mathcal{L}_{\text{adv}}(D(\mathcal{E}(x)))$
$+\log(D_{\psi}(x))$
$+\mathcal{L}_{\text{reg}}(x;\mathcal{E},D)\big{)}$
where $\mathcal{L}_{rec}$ is a combination of L1, perceptual losses for the RGB channels, and Pearson Correlation Coefficient (PCC), TV regularization losses for the depth channels. $\mathcal{L}_{\text{adv}}$ is the adversarial loss, $D_{\psi}$ is a patch-based discriminator loss, and $\mathcal{L}_{\text{reg}}$ is the KL-regularisation loss. The incorporation of PCC and TV terms for the depth channels leads to better retention of high-frequency details in the reconstructed depth map, as observed in Fig 4. We finetune this VAE on a subset of our dataset for 5000 training steps with batch size 16 and learning rate 1e-05.
3.3.3 UNet Finetuning
With our finetuned autoencoder, we next train the UNet with the frozen VAE on $\mathcal{X}$ with the following objective:
$\mathcal{L}=\mathbb{E}_{i\sim\mathcal{U}(N),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\Big{[}$ $\|\epsilon_{t}-\bm{\epsilon_{\theta}}(\mathbf{z}^{i}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},c_{attn},c_{\theta}(y^{i}))\|_{2}^{2}\Big{]}$ (6)
Input RGB Recon. RGB Input Depth Recon. Depth
Ours (finetuned) Refer to caption Refer to caption Refer to caption Refer to caption
LDM3D VAE Refer to caption Refer to caption Refer to caption Refer to caption
Ours (finetuned) Refer to caption Refer to caption Refer to caption Refer to caption
LDM3D VAE Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: RGBD reconstruction comparison of our finetuned VAE with the LDM3D VAE (Stan et al., 2023). Unlike LDM3D, our VAE finetuned on a synthetic dataset preserves sharp details and edges of the input depth map while also preventing color artifacts in RGB.
where $\mathbf{z}_{t}^{i}$ is the encoded image ${I}^{i}$ diffused with sampled noise $\epsilon$ at time step $t$. Thus, the UNet learns to generate clean RGBD images $I^{i}$, conditioned on artifact RGBD images $I^{i}$, confidence maps $\mathcal{C}^{i}$, CLIP features $c_{attn}$, and instructions $y^{i}$.
3.3.4 Inference using RGBD Novel View Synthesis
At inference time, given a render and depth map with artifacts and the base instruction $y=$“Inpaint low-confidence regions in the image and depth map with coherent details and remove all floaters and Gaussian artifacts." the finetuned UNet $\bm{\epsilon_{\theta}}$ learns to predict the noise in latent $\mathbf{z}_{t}$ according to $t\sim\mathcal{U}[t_{min},t_{max}]$ as:
$\hat{\epsilon}_{t}=$ $\,\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},c_{attn},\varnothing)+s_{I}(\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},c_{attn},\varnothing)-\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},c_{attn},\varnothing))$ (7)
$+s_{T}(\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\mathcal{\hat{C}},c_{attn},c_{\theta}(y))-\bm{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\hat{I}),\mathcal{\hat{C}},c_{attn},\varnothing))$
where $s_{I}$ and $s_{T}$ are the RGBD image and prompt guidance scales, dictating how strongly the final multistep reconstruction agrees with the RGBD render $\hat{I}$ and the edit prompt $y$, respectively. After $k=20$ DDIM Song et al. (2021) sampling steps, we obtain our final RGBD render by decoding the denoised latent as $x_{\pi}=[I_{\pi},D_{\pi}]=\mathcal{D}(z_{0})$.
3.4 Scene Reconstruction with Diffusion Priors
Our diffusion priors infer plausible detail in unobserved regions. However, the generated images at novel poses lack 3D consistency due to lack of explicit 3D conditioning of the UNet through pose embeddings. For this, we devise an iterative strategy where we first sample novel views along a B-spline trajectory fitted to the training views. We initialize the Gaussian optimization with the set of Gaussians $\mathcal{G}$ fitted to the training views. Each novel view is added to the training stack at the beginning of every densification step to encourage the optimization to adjust to the distilled scene priors. At every iteration, we sample either an observed or unobserved viewpoint from the current training stack. We jointly optimize Gaussian attributes, and both train and novel poses. We employ the 3DGS objective for the training views. For novel views, we employ the SparseFusion Zhou and Tulsiani (2023) objective in the RGB space (Wu et al., 2024) and a PCC loss for the rendered and denoised depths.
$\mathcal{L}_{sample}(\psi)$ $=\mathbb{E}_{\pi,t}\Big{[}w(t)(\|I_{\pi}-\hat{I}_{\pi}\|_{1}+\mathcal{L}_{p}(I_{\pi},\hat{I}_{\pi}))\Big{]}+w_{d}\cdot PCC(D_{\pi},\hat{D}_{\pi})$ (8)
where $\mathcal{L}_{p}$ is the perceptual loss Zhang et al. (2018), $w(t)$ a noise-dependent weighting function, $I_{\pi},D_{\pi}$ are the rendered image and depth at novel viewpoint $\pi$, and $\hat{I}_{\pi},\hat{D}_{\pi}$ are their rectified versions obtained with our diffusion prior.
3.5 Test-time Pose Alignment
GScenes reconstructs a plausible 3D scene from unposed source images. However, reconstruction with few views is inherently ambiguous as several solutions can satisfy the train view constraints. Hence, the reconstructed scene would most likely be quite different from the actual scene from which $M$ views were sampled. Hence, for a given set of test views, following prior work Fan et al. (2024); Jiang et al. (2024), we freeze the Gaussian attributes and optimize the camera pose for each target view by minimizing a photometric loss between the rendered image and test view. Following this alignment step performed for $500$ iterations per test image, we evaluate the NVS quality.
InstantSplat RegNeRF DiffusioNeRF ZeroNVS* COGS Ours Ground Truth
Stump(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kitchen(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Bicycle(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Garden(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Treehill(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of GScenes with few-view methods. Our approach consistently fairs better in recovering image structure from foggy geometry, where baselines typically struggle with “floaters” and color artifacts.
4 Experiments
We compare GScenes with state-of-the-art pose-free and pose-required sparse-view reconstruction methods in Fig and Table . We also ablate the different components and design choices of our diffusion model.
4.1 Experimental Setup
Evaluation Dataset
We evaluate GScenes on the 9 scenes of the MipNeRF360 dataset (Barron et al., 2022), comprising 5 outdoor and 4 indoor scenes. Each scene has a central object or area of complex geometry with an equally intricate background. This makes it the most challenging $360^{\circ}$ dataset compared to CO3D (Reizenstein et al., 2021), RealEstate 10K (Zhou et al., 2018), DTU (Jensen et al., 2014), etc. We pick the $M$-view splits as proposed by ReconFusion Wu et al. (2024) and CAT3D Gao* et al. (2024) and evaluate all baselines on the official test views where every $8^{th}$ image is held out for testing.
Fine-tuning Dataset
We fine-tune our diffusion model on a mix of 165 scenes encompassing Tanks and Temples (Knapitsch et al., 2017), CO3D (Reizenstein et al., 2021), DTU (Jensen et al., 2014), Deep Blending (Hedman et al., 2018), and LLFF (Mildenhall et al., 2019) to obtain a total of $28161$ data samples. We first train 3DGS on sparse and dense subsets of each scene for $M\in\{3,6,9,18\}$. For $M>18$, novel view renders and depth maps mostly show Gaussian blur as artifacts. Finetuning this model takes about $1$-day on a single A6000 GPU.
Metrics
Our quantitative metrics are used to evaluate two tasks - quality of novel views post reconstruction and camera pose estimation. For the former, we compute 3 groups of metrics - FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) due to the generative nature of our approach, perceptual metrics LPIPS (Zhang et al., 2018) and DISTS (Ding et al., 2020) to measure similarity in image structure and texture in the feature space, and pixel-aligned metrics PSNR and SSIM. However, PSNR and SSIM are not suitable evaluators of generative techniques (Chan et al., 2023; Sargent et al., 2024) as they favor pixel-aligned blurry estimates over high-frequency details.
Baselines
We compare our approach against $8$ baselines. FreeNeRF (Yang et al., 2023), RegNeRF (Niemeyer et al., 2022), DiffusioNeRF (Wynn and Turmukhambetov, 2023) are pose-required few-view regularization methods based on NeRFs. ZeroNVS (Sargent et al., 2024) reconstructs a complete 3D scene from a single image using a novel camera normalization scheme and anchored SDS loss. We use the ZeroNVS baseline introduced in ReconFusion (Wu et al., 2024), designed to adapt ZeroNVS to multi-view inputs. ReconFusion and CAT3D are state-of-the-art multi-view conditioned diffusion models for sparse-view reconstruction. We use the reported average performance on the $9$ scenes of MipNeRF360 with classical metrics - PSNR, SSIM, and LPIPS for quantitative comparison and pick the relevant scenes and test views from the two papers for qualitative comparison. For pose-free methods, we pick InstantSplat (Fan et al., 2024) and COGS (Jiang et al., 2024) and show that our approach outperforms both in NVS qualitatively and quantitatively.
4.2 Implementation Details
We implement our entire framework in PyTorch 2.3.1 and run all experiments on single A5000 or A6000 GPUs. We render images and depth maps in the 400-600 pixel range as this is closest to the output resolution of Stable Diffusion. We finetuned our diffusion model for $10k$ iterations with batch size $16$ and learning rate 1e-4. To enable classifier-free guidance, we randomly dropout conditioning inputs like text, image, confidence map, etc., with probability of $0.05$ during training.
We fit 3D Gaussians to sparse input images and a DUSt3r point cloud and train it for $1k$ iterations as in  Fan et al. (2024) to obtain our initial Gaussian representation $\mathcal{G}$. The classifier-free guidance scales are set to $s_{I}=2.5$ and $s_{T}=7.5$. We use $t_{max}=0.99$ and linearly decrease $t_{min}$ from $0.98$ to $0.70$. We sample novel-view renders and depth maps for $k=20$ DDIM sampling steps. $w_{d}$ is decayed linearly from $1$ to $0.01$. We also linearly decay the weight of $\mathcal{L}_{sample}$ from $1$ to $0.1$ over $10k$ iterations. GScenes reconstructs the complete 3D scene in about $5$ minutes on a single A6000 GPU.
4.3 Comparison Results
Table 1: Quantitative comparison with state-of-the-art sparse-view reconstruction techniques on classical metrics.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Method 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
FreeNeRF 11.888 12.877 13.680 0.146 0.180 0.197 0.675 0.654 0.638
RegNeRF 12.297 13.209 13.802 0.147 0.170 0.180 0.668 0.656 0.625
DiffusioNeRF 13.134 16.191 16.732 0.167 0.283 0.337 0.680 0.543 0.530
ZeroNVS* 11.902 11.789 11.729 0.142 0.133 0.128 0.710 0.702 0.694
ReconFusion 15.50 16.93 18.19 0.358 0.401 0.432 0.585 0.544 0.511
CAT3D 16.62 17.72 18.67 0.377 0.425 0.460 0.515 0.482 0.460
InstantSplat 13.657 14.733 14.926 0.222 0.265 0.246 0.603 0.540 0.590
COGS 12.267 12.428 13.030 0.173 0.186 0.201 0.616 0.632 0.596
GScenes 14.976 15.972 16.563 0.284 0.295 0.314 0.593 0.517 0.577
Table 2: Quantitative comparison with few-view reconstruction techniques on metrics suited for generative reconstruction.
FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$
Method 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
FreeNeRF 342.484 343.378 334.805 0.252 0.262 0.268 0.376 0.363 0.357
RegNeRF 343.593 335.606 325.228 0.260 0.276 0.257 0.387 0.383 0.371
DiffusioNeRF 323.047 269.995 246.265 0.230 0.166 0.141 0.395 0.325 0.328
ZeroNVS* 356.395 350.362 343.930 0.283 0.294 0.299 0.433 0.427 0.413
InstantSplat 268.526 226.110 242.853 0.191 0.131 0.140 0.294 0.298 0.310
COGS 231.274 274.827 248.220 0.136 0.183 0.152 0.288 0.310 0.286
GScenes 253.659 220.292 210.255 0.176 0.098 0.106 0.284 0.245 0.241
We report qualitative and quantitative comparisons of GScenes against all related baselines in Fig 5 and Tables 1 and  2. Out of the $8$ baselines, only InstantSplat and COGS are pose-free techniques, while the remaining require ground truth poses for both training and evaluation. Note that COGS relies on ground truth camera intrinsics while GScenes and InstantSplat do not. On the classical metrics (Tab 1), we are third behind the 2 SOTA posed reconstruction methods - CAT3D and ReconFusion. Except for FID and KID on the 3-view split, we outperform all related baselines in Table 2. Due to the unavailability of open-source code, evaluating ReconFusion and CAT3D on these measures is unfortunately not possible.
Base w/o conf. map w/o instruction w/o depth Ours Ground Truth
Counter(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Treehill(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Treehill(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Ablation Study with the Counter and Treehill scenes of MipNeRF360. Images above are heldout test-view renders obtained after test-time pose alignment (Sec 3.5). All of our proposed conditionings improve the diffusion model’s ability to inpaint missing details and remove artifacts in novel view renders.
4.4 Ablation Studies
In Fig 6, we ablate different components of our diffusion model. We pick the counter and treehill scenes for this experiment. The Base variant only performs image-to-image diffusion in the RGB space and tries to fill in missing details and remove artifacts based on the partial information available from the novel view render. Conditioning with our pixel-aligned confidence measure leads to better detection and elimination of Gaussian artifacts. Foreground and background artifacts in both the render and depth map are not effectively dealt with in the absence of this conditioning input. Performing diffusion only in the RGB space usually produces depth maps with sharp foreground artifacts. Incorporating joint modeling of RGB and depth through our RGBD diffusion model and subsequent depth optimization during scene reconstruction (Sec 3.4) produces smoother depth maps while also restricting floaters and Gaussian artifacts. As we randomly dropout conditioning prompts (instruction, confidence map, etc.) during fine-tuning, the diffusion model is adept at image synthesis even in the absence of the base instruction prompt. However, an intuitive instruction like “Inpaint low-confidence regions in the image and depth map with coherent details and remove all floaters and Gaussian artifacts.” usually guides the model better in localizing empty regions and Gaussian artifacts.
Table 3: Pose estimation accuracy with GScenes. SfM-poses for training views estimated from the full observation set are used as ground truth. We report errors in camera rotation and translation using Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) as in  Bian et al. (2023). Despite the DUSt3r initialization of camera extrinsics and subsequent pose optimization, there are large errors w.r.t COLMAP poses of the dense observation set, harming test-pose alignment and subsequent NVS quality.
RPEt $\downarrow$ RPEr $\downarrow$ ATE $\downarrow$
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
GScenes 37.684 27.893 30.916 124.501 58.915 40.424 0.261 0.294 0.266
4.5 Limitations & Future Work
GScenes is a first step towards a generative solution for pose-free sparse-view reconstruction of large complex scenes. However, it is not free of its fair share of limitations. The quality of the final reconstruction depends heavily on the initial relative pose estimation by the DUSt3r pipeline, and even though the poses are further optimized jointly with Gaussian attributes during training, there are still large differences with the ground truth COLMAP poses estimated from dense views as we show in Tab 3. This limits fair comparison with posed reconstruction methods, as even the test-time pose alignment step (Sec 3.5) cannot compensate for the initial errors in the pipeline. Finally, our diffusion model is limited in its 3D outpainting abilities as, unlike a simplified 2-view setting, there is no explicit view conditioning possible in a general pose-free N-view setting. Existing state-of-the-art posed reconstruction techniques like ReconFusion and CAT3D are able to leverage stronger diffusion priors for scene reconstruction as the diffusion models have view conditioning in the form of pixelNeRF embeddings or “raymap” representations. We hope to address some of these limitations as part of our future work.
5 Conclusion
In this work, we present GScenes where we integrate an instruction-following RGBD diffusion model with a pose-free reconstruction pipeline in InstantSplat to reconstruct a $360^{\circ}$ 3D scene from a few uncalibrated 2D images. We also introduce a pixel-aligned confidence measure to further guide the diffusion model in uncertain regions with missing details and artifacts. Our experiments show that GScenes outperforms existing pose-free reconstruction methods in scene reconstruction and performs competitively with state-of-the-art posed sparse-view reconstruction methods.